Consider the initial value problem y + y – 12y = 0, y(0) = a, y(0) = 5 Find the value of a so that the solution to the in
Answer #1

Solution : y+y-12y = 0, 7(0) = d, y(o)=5 : Characteristic equation of given. De is, m2 + m -12 = 0 - m2 +47-3m-12=0 : mcm+